OpenStudy (anonymous):

what is 4(3-y)=6-2(1-3y) ?

OpenStudy (anonymous):

Do you know how to expand brackets?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Distribute the 4 and the 2 on each side.

OpenStudy (anonymous):

It's called the distributive property: a(b+c)=ab+ac So in this case 4(3-y)=4*3-3*y. Can you do it for 2(1-3y)?

OpenStudy (anonymous):